OpenStudy (anonymous):

Hi, I'm having a bit of trouble finding the integral for 1/(1+x^7) and e^(-x^2) could someone point me in the right direction please?

OpenStudy (bahrom7893):

try wolframalpha.com it shows steps too

OpenStudy (anonymous):

not for these two, the steps are unavailable. If you can see the procedures to take let me know

OpenStudy (amistre64):

e^(-x^2) isnt this the same as e^(-2x) ?

OpenStudy (amistre64):

it aint :)

OpenStudy (anonymous):

I don't think so,

OpenStudy (amistre64):

there are plenty of functions that are simply not integrateable

OpenStudy (anonymous):

true, so do you see any integration technique available to a Calculus 2 student that would solve these 2 equations?

OpenStudy (amistre64):

i got books in the library here with tables of integration ... much more extensive than the leaflets in the textbook

OpenStudy (anonymous):

would it not be ln[1+x^7]+C?

OpenStudy (amistre64):

take the derivative to test it out :)

OpenStudy (amistre64):

there is a 7x^6 missing for that to be good; maybe try int by parts?

OpenStudy (amistre64):

or fancy substitutions ;)

OpenStudy (anonymous):

hmm, ok, ill see about those

OpenStudy (amistre64):

suppose we make: x = z^(1/7)

OpenStudy (amistre64):

dx = (z^6)/7 dz \[\int\frac{1}{1+(x)^7}dx\] \[\int\frac{z^6}{7(1+(z^{1/7})^7)}dz\]

OpenStudy (amistre64):

i messed up the derivative of z^(1/7) i think :)

OpenStudy (anonymous):

-6/7

OpenStudy (amistre64):

im gonna switch to another variable that stands out better \[[a^{1/7}]'=\frac{1}{7a^{6/7}}\] \[dx = \frac{1}{7a^{6/7}}dz\]

OpenStudy (amistre64):

\[\int\frac{1}{7a^{6/7}(1+a)}da\] is that workable?

OpenStudy (anonymous):

I think we could int by parts now

OpenStudy (anonymous):

I don't know, I've never done sometihng like this but I have a=7a^6/7, could this lead to an answer?

OpenStudy (amistre64):

cant really say fer sure. Ive never tried it out b4 either :) still leafing thru some pages tho

OpenStudy (anonymous):

I think the questions are bogus, e^(-x/2) involves an error function which we havent learned yet, and the answer for 1/(1+x^7) is humongus

OpenStudy (anonymous):

in other words, I think we are supposed to realize that these are unsolvable for us

OpenStudy (anonymous):

thx for the help :)

OpenStudy (amistre64):

yep; that right :) my text says: "you should realize that no table is adequate for all integration problems. A simple example is \(\int\ e^{-x^2}\). No elementary function exists whose derivative is \(e^{-x^2}\); in such a case it is necessary to restort to approx. techniques.